OpenStudy (anonymous):

An escalater is 65m long and inclined 30 degrees. Determine the work done by the escalaters step acting on a 57kg person that rides this distance.

OpenStudy (michele_laino):

hint:  the force acting on the person has to be equal to the parallel component of the weight of that person with respect to the ascalater. Now that force has the subsequent magnitude: \[\Large F = mg\sin 30\] where g is the gravity, namely g=9.81 m/sec^2 So the requested work W is: \[\Large W = F \times d = mgd\sin 30\] where d= 65 meters

OpenStudy (michele_laino):

oops..escalator*